OpenStudy (anonymous):

Which of the following is an example of conditional probability? A. Probability of getting an A on a test B. Probability of getting an A on a test, given that you didn't fail the test C. Probability of drawing a club from a deck of 52 cards D. Probability of hitting a home run

OpenStudy (anonymous):

@FortyTheRapper

OpenStudy (fortytherapper):

A conditional probability is basically something like... This could happen, as long as this did. There's only one answer that has that type of structure. Know which one?

OpenStudy (anonymous):

B"

OpenStudy (anonymous):

?

OpenStudy (fortytherapper):

Exactly

OpenStudy (anonymous):

thanks man. u always come through

OpenStudy (anonymous):

u should be a qualified helper

OpenStudy (fortytherapper):

No problemo

OpenStudy (fortytherapper):

Maybe one day. This is just my second day here. It's addicting to me lol

OpenStudy (anonymous):

do you have time for a couple more?

OpenStudy (fortytherapper):

Sure

OpenStudy (anonymous):

can be read as "the probability that A occurs given that B has occurred." A. True B. False

OpenStudy (anonymous):

P(A|B)

OpenStudy (anonymous):

is in the beginning

OpenStudy (fortytherapper):

Yes. It's true if this is still about conditional probablities

OpenStudy (anonymous):

ye that one was ez

OpenStudy (anonymous):

u got soundcloud?

OpenStudy (fortytherapper):

Yeah, I haven't posted anything there for like 6 months (Nothing serious for 8 months)

OpenStudy (anonymous):

damn, u dont want this work in a rap battle

OpenStudy (anonymous):

What is the probability of rolling a number greater than or equal to 8 with two dice, given that at least one of the dice must show a 6? 9/11 2/11 7/11 4/11

OpenStudy (anonymous):

9/11?

OpenStudy (fortytherapper):

Hmm, I don't know about this one. Since 1 must show a 6, that makes the probability 1/6. 8-6 = 2 That means the second dice must show a 2 or greater in order to meet this requirement. 2,3,4,5,6... that's 5 numbers, so 5/6 for the second dice. But, that doesn't match

OpenStudy (anonymous):

yea this website my school uses is dumb

OpenStudy (anonymous):

If there is a 90% chance of rain tomorrow and a 70% chance of wind and rain, what is the probability that it is windy, given that it is rainy? Round your answer to the nearest percent. A. 78% B. 33% C. 71% D. 67%

OpenStudy (anonymous):

D?

OpenStudy (fortytherapper):

Yeah, I never took Statistics, so I'm limited there I would convert those numbers to equal out of 1, not 100. So you would get .9 and .7. Then divide those two numbers (.7/.9) Remember to round

OpenStudy (anonymous):

1 more please?

OpenStudy (anonymous):

According to the general equation for conditional probability, if p (a^b')=1/6 and p(b')=7/24 , what is p(a[b) ? 3/7 4/7 5/7 2/7

OpenStudy (anonymous):

2/7?